OpenStudy (anonymous):

evaluate the expression and write the result in the form a + bi. (5 + 3i)(2 − i) over/3 + i

OpenStudy (anonymous):

10 - 5i + 6i -3i^2 = 10 + i + 3 = 13 + i.

OpenStudy (anonymous):

:^)

OpenStudy (lalaly):

(5+3i)(2-i)= 10-5i+6i+3 = 13 + i (13+i)/ (3+i)= 4-i

OpenStudy (anonymous):

so wait which one is it?

OpenStudy (lalaly):

4- i

OpenStudy (anonymous):

I got the numerator. The whole problem is now reduced to (13 + i)/(3 + i).

OpenStudy (anonymous):

= (13 + i)(3 - i)/[3 + i)(3 = i)]

OpenStudy (anonymous):

= (39 - 13i + 3i - i^2)/10 (40 - 10i)/10 4 - i.